HIGHLY EXCITED STATES OF Cs ATOMS ON HELIUM NANODROPLETS

\underline{F. LACKNER}, M. THEISEN, M. KOCH, and W.E. ERNST, Institute of Experimental Physics, Graz University of Technology, Petersgasse 16, A-8010 Graz, Austria.

Cs atoms on the surface of helium nanodroplets have been excited to high lying nS (n = 8-11), nP (n = 8-11), and nD (n = 6-10) levels. A two-step excitation scheme^{a} via the 6^{2}P_{1/2}(^{2}Π_{1/2}) state using two cw lasers was applied. This intermediate state has the advantage that a large fraction of the excited Cs atoms does not desorb from the helium nanodroplets. An absorption spectrum was recorded by detecting laser induced fluorescence light from the 6^{2}P_{3/2} →6^{2}S_{1/2} transition. The pseudo-diatomic model for helium nanodroplets doped with single alkali-metal atoms holds for the observed spectrum. An investigation of spectral trends shows that the n’^{2}P(Π)←6^{2}P_{1/2}(^{2}Π_{1/2}) and n’^{2}D(Δ)←6^{2}P_{1/2}(^{2}Π_{1/2}) (n’ > 9) transitions are lower in energy than the corresponding free-atom transitions. This indicates that the Cs*–He_{N} potential becomes attractive for these highly excited states. Our results suggest a possibility of generating an artificial super-atom with a positive ion core inside a helium nanodroplet and the electron outside, which will be subject to future experiments.

aM. Theisen, F. Lackner, F. Ancilotto, C. Callegari, and W.E. Ernst, Eur. Phys. J. D 61, 403-408 (2011)